CHEMICAL MODELING Of O_{2} IN INTERSTELLAR CLOUDS

\underline{DONGHUI QUAN}, Chemical Physics Program, The Ohio State University, Columbus, OH 43210; ERIC HERBST, Departments of Physics, Astronomy, & Chemistry, The Ohio State University, Columbus, OH 43210; GEORGE E. HASSEL, Department of Physics, The Ohio State University, Columbus, OH 43210.

In cold cores of interstellar clouds, molecular oxygen cannot be detected at all down to rather low upper limits. This has been a problem for chemistry modelers since steady-state results of their models always give high O_{2} abundances. In chemical models, the reaction between O and OH is the main source of O_{2} formation. Experimental studies have shown that the rate coefficient of this reaction down to 39 K is 3.5 × 10^{−11} cm^{3} s^{−1}, which is almost one order of magnitude lower than values used by modelers. Two recent quantum calculations with an accurate HO_{2} potential surface have suggested that at 10 K, this rate coefficient is even much smaller. This small rate coefficient might inhibit the production of O_{2} and explain the negative results for O_{2} towards cold interstellar clouds. In this work, we show how the interstellar O_{2} abundance is affected when the rate coefficient is decreased. Under standard O-rich elemental abundances, the calculated O_{2} abundance is sufficiently low to lie below the observed upper limit only at early times with all the rate coefficient values we investigated. Under C-rich abundances, both early-time and late-time solutions are possible.